1. One hundred unit point masses are connected with identical linear springs of unit stiffness

on a circular ring of unit radius. Find the largest angular frequency of oscillations of the masses.


  1. One mole of argon (assumed to be an ideal gas) expands according to the relation p =

kV between its pressure and volume,k = const. Find the entropy change of the gas when its volume

doubles.


1
Expert's answer
2021-05-12T17:55:38-0400

1.

"\omega=\sqrt \frac km=\sqrt \frac{nk_0}{m}=\sqrt \frac{100\cdot 1}{1}=10."

2.

"\Delta S=\nu R\ln \frac{V_2}{V_1}=1\cdot 8.31\cdot 0.69=5.76~\frac{J}{K}."